Use the following information: A container has 8360.0 torr of Nitrogen, 5.00 atm of Oxygen, 101.1 kPa of Argon, and 760.0 mm Hg of Carbon Dioxide.


What is the mole fraction of Carbon Dioxide?


A. 0.0556

B. 0.0670

C. 0.2567

D. 0.6676

E. 0.7572


1
Expert's answer
2021-05-12T06:30:20-0400

"\frac {8360 torr × 1 atm} {760 torr}" = 11 atm N2

5 atm O2


"\frac {101.1 kPa × 1 atm} {101.32}" = 0.998 atm Ar


"\frac {760 mm Hg × 1 atm} {760 mm Hg}" = 1 atm CO2


Total Pressure = 11 atm + 5 atm + 0.998 atm +1 atm = 17.998 atm


mole fraction = "\frac {5 atm} {17.998}"


Mole fraction of Carbon Dioxide= 0.278